OpenStudy (anonymous):

the mathematical model w/2-3h=53 describes a weight, w, in pounds, that lies within the healthy weight range for a person whose height is h inches over 5 feet. use the formula to find a healthy weight for a person whose height is 5'11".

OpenStudy (shane_b):

Is that \[\frac{w}{2}-3h=53\]?

OpenStudy (anonymous):

yes

OpenStudy (shane_b):

Do you know how to solve that equation for w?

OpenStudy (anonymous):

i thought i did but i'm getting it wrong

OpenStudy (shane_b):

Ok no problem. What's the first step that you'd use to solve it for w?

OpenStudy (anonymous):

multiply the equation by 2 right?

OpenStudy (shane_b):

You could do that...or you could also just add 3h to both sides. Pick one and show me what you get

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

w-30=106

OpenStudy (anonymous):

add 3o to both sides

OpenStudy (shane_b):

Whoa :)

OpenStudy (anonymous):

w=136

OpenStudy (shane_b):

Ok...let me show you how to do it. Which method would you like to start with?

OpenStudy (anonymous):

3h to both sides

OpenStudy (shane_b):

ok. \[\frac{w}{2}-3h=53\]\[\frac{w}{2}\cancel{-3h}+\cancel{3h}=53+3h\]\[\frac{w}{2}=53+3h\] Now you need to multiply everything by 2:\[(\cancel {2})\frac{w}{\cancel{2}}=2(53+3h)=2*53+2*3h\]Now reduce: \[w=106+6h\]

OpenStudy (shane_b):

The only thing left is to convert 5'11" to inches and plug it in for the height and solve.

OpenStudy (anonymous):

oh okay hehehe

OpenStudy (anonymous):

thanks

OpenStudy (shane_b):

yw :)